Three major contributing resonance structures are possible for the following cation. One is given below. Draw the remaining structures (in any order), including nonbonding electrons and formal charges

Three major contributing resonance structures are
Answer #1
Concepts and reason

Resonance is a method of representing the delocalized electrons by Lewis structures of a molecule. Resonance structure is stabilized by the delocalization of π-bond.

Resonance structures are drawn by denoting the electron flow to through the more electronegative atom. In the resonating structure unsaturated carbon is present that mean sp2 – hybridized orbital.

Formal charge: Is the method of representing the number of electron associated with every atom in the molecule, it is much useful in identifying the reactive site in a molecule.

Fundamentals

Flow of electrons in resonance:
avon
electron transforms from a bond to lone pair
electron transforms from lone pair to a bon

In the resonance structure charge are always same. Position of atoms should not be changed. Only charge transfer is taking place.

General procedure for drawing resonance structures:

1. Movement of electron always takes place from electronically rich center to electron deficient center.

2. Number of charge should be same.

3. Migration of double bond takes place within adjacent atom.

4. Position of the atoms should be same.

5. Identify the formal charge.

The formal charge on an atom could be calculated as followed:

Formalcharge(FC)=(numberofvalenceelectronsintheneutralatom)(numberofelectronsinbonding)2(numberofelectronsaslonepair){\rm{Formal}}\,{\rm{charge (F}}\,{\rm{C)}}\,{\rm{ = }}\,\left( \begin{array}{l}\\{\rm{number}}\,{\rm{of}}\,{\rm{valence}}\,{\rm{electrons}}\,\\\\{\rm{in}}\,{\rm{the}}\,{\rm{neutral}}\,{\rm{atom}}\\\end{array} \right)\,{\rm{ - }}\frac{{\left( \begin{array}{l}\\{\rm{number}}\,{\rm{of}}\,{\rm{electrons}}\,\\\\{\rm{in}}\,{\rm{bonding}}\\\end{array} \right)}}{{\rm{2}}}{\rm{ - }}\left( \begin{array}{l}\\{\rm{number}}\,{\rm{of}}\,{\rm{electrons}}\,\\\\{\rm{as}}\,\,{\rm{lone}}\,{\rm{pair}}\\\end{array} \right)

-T
A
པ། བབ། །
CH3
nucleophilic center
electrophilic center
electrophilic center

Formal charge is assigned according to the movement of electrons. Formal charge of each atom is +1

-I
CHO
is
-
ylpy
H
19
I-

Ans:

Resonance structure for the given carbocation is depicted below:

T
لای
و
H
لای
CH3
-CH₃